OpenStudy (anonymous):

-5/12 * -9 * (-3/20)

OpenStudy (anonymous):

-(9/16)

OpenStudy (anonymous):

sorry, -9/16

OpenStudy (anonymous):

\[-\frac{5}{12}\times -\frac{9}{1} \times -\frac{3}{20}=(-1)(-1)(-1)\frac{5 \times 9 \times 3}{12 \times 1 \times 20}=-\frac{135}{240}=?\]

jhonyy9 (jhonyy9):

-5/12 *-9* -3/20 -5/4 *-3* -3/20 -1/4 *-3* -3/4 -9/16